Main page‎ > ‎Observing Runs‎ > ‎Night Reports‎ > ‎

Night report page - 20130726


Observers: Paul Vreeswijk, Annalisa De Cia and Cristina Barbarino
Support team: Ofer Yaron and Avishay Gal-Yam
TAT on duty: Maria Teresa Botticella and Matt Nicholl
TIO: Duncan Castex

************************************************

Subrun.night: 1.1
Instruments: EFOSC2

*** WEATHER CONDITIONS

Night  Humidity  Wind speed      Sky
 
beg.    ~35%       15-20 m/s     clear
mid.    ~35%       17-22 m/s     some cirrus
end.    ~35%       15-20 m/s     some cirrus


*** AFTERNOON CALIBRATIONS
 
PESSTO_calib
PESSTO_bias


*** LOG

Dome closed all night due to high NE wind. NO OBSERVATIONS.


*** NOTES

* SN in M74: http://www.cbat.eps.harvard.edu/unconf/followups/J01560719+1738468.html
* Internal flats for grisms #13 and #16 show strong (water?) absorption around 7000 AA (and a weaker one around 9000 AA for grism #16)
Comments